What is the solution for homogeneous equation (2x+y)2dx=xydy


1
Expert's answer
2021-02-23T11:25:44-0500

What is the solution for homogeneous equation (2x+y)2dx=xydy

Solution:

"y=ux" , "dy=udx+xdu" .

"(2x+ux)^2dx=x^2u(udx+xdu)"

"4x^2(1+u)dx=ux^3du"

"4\displaystyle\frac{dx}{x}=\frac{u}{1+u}du"

"4\displaystyle\frac{dx}{x}=(1-\frac{1}{1+u})du"

"\ln{C}+4\ln{|x|}=u-\ln{|1+u|}"

"Cx^4=\displaystyle\frac{e^u}{1+u}"

"Cx^4=\displaystyle\frac{e^\frac yx}{1+\frac yx}"

"Cx^5=\displaystyle\frac{e^\frac yx}{x+y}"

Answer: "Cx^5=\displaystyle\frac{e^\frac yx}{x+y}" .